The square of Mark's age 3 years ago is 6 times the age he will be in 9 years. What is his age now?
Answer by texttutoring(324)

The square of Mark's age 3 years ago

Let x = Mark's age now.
3 years ago, he was (x-3) years old. The square of his age 3 years ago is (x-3)^2


In 9 years (x+9), Mark will be 6 times the square of his (x-3) age.

(x-3)^2 = 6(x+9)
(x^2-6x+9) = 6x+54
x^2 - 12x -45 = 0

Factor this. Think what numbers multiply to -45 but add up to -12? -15 and 3.

(x-15)(x+3)=0
x=15, or x=-3

We can eliminate x=-3, because Mark can't have a negative age.
Mark is 15 years old now.